# The Axiom of Real Determinacy and the Axiom of Real Blackwell Determinacy

### v1

@@ Abstract.
The research in this paper was partially conducted when the authors stayed in the Mittag-Leffler Institute for the semester program “Mathematical Logic: Set Theory and Model Theory” in fall 2009. The authors are grateful to the Mittag-Leffler Institute for their hospitality and support. This paper contains results from the first author’s Ph.D. thesis [6]. The first author would like to thank Benedikt Löwe for his supervision. The first author’s research is partially supported by the Guangdong Philosophy and Social Science Foundation with grant number GD25YZX01.
@@ 1. Introduction
In 1998, Martin [16] proved that in most cases, Blackwell determinacy axioms follow from the corresponding determinacy axiom. Martin conjectured that they are equivalent, and many instances of equivalence have been shown (e.g., [17] and Martin’s proof of $\mathbf{\Pi}^{1}_{1}$ determinacy presented in [14, Corollary 3.9]). However, the general question, and in particular the most intriguing instance, namely whether $\mathsf{AD}$ and the axiom of Blackwell determinacy $\mathsf{Bl}\text{-}\mathsf{AD}$ are equivalent, remain open.
@@
Solovay [20] proved that the consistency of $\mathsf{ZF}$$\mathsf{DC}$$\mathsf{AD}_{\mathbb{R}}$ is strinctly stronger than that of $\mathsf{ZF}$$\mathsf{AD}_{\mathbb{R}}$. Hence assuming $\mathsf{DC}$ is not optimal when considering the equivalence between $\mathsf{AD}_{\mathbb{R}}$ and $\mathsf{Bl}\text{-}\mathsf{AD}_{\mathbb{R}}$. While we do not know whether they are equivalent in $\mathsf{ZF}$$\mathsf{AC}_{\omega}(\mathbb{R})$, we show that they are equiconsistent:
@@ 2.1. Measure and category
Suppose that every set of reals is Lebesgue measurable. Then for every Borel probability $\mu$ on the Cantor space, every set of reals is $\mu$-measurable and the ideal of $\mu$-null sets is closed under well-ordered unions. Also, there is no injection from $\omega_{1}$ to the reals.
@@ 2.3. Blackwell games
For the arguments for the assertion that $\Gamma$ is a boldface pointclass as well as the closure of $\Gamma$ under countable unions and intersections, and $\exists^{\mathbb{R}}$, see [18, 2B.2]. For the closure of $\Gamma$ under $\forall^{\mathbb{R}}$, let $\Delta$ be the pointclass of all the Susloin and co-Suslin sets. Then by $\mathsf{Bl}\text{-}\mathsf{AD}$ and Theorem 13, every set of reals in $\Delta$ is determined. Now by the second periodicity theorem (see [18, 6C.3]), $\Gamma$ is closed under $\forall^{\mathbb{R}}$. ∎
@@ 2.6. Wadge games and Blackwell determinacy
where $\mu_{x,\tau}=\mu_{\sigma_{x},\tau}$ and $\sigma_{x}$ is the mixed strategy for player I concentrating on $x$ of measure one as in § 2.3. It is easy to see that the right hand side of the equivalence is $\undertilde{\mathbf{\Sigma}}^{1}_{2}$ in $B$. If player I has an optimal strategy in $\mathcal{G}_{\mathrm{W}}(A,B)$, then one can prove that $B$ is $\undertilde{\mathbf{\Sigma}}^{1}_{2}$ in $A^{\mathrm{c}}(=2^{\omega}\setminus A)$ in the same way and hence $B$ is $\undertilde{\mathbf{\Sigma}}^{1}_{2}$ in $A$. ∎
@@ 2.7. A weak version of Moschovakis’ Coding Lemma
Recall from § 2.3 that for a real $b$, $\tau_{b}$ is the mixed strategy for player II concentrating on $b$ of measure one. Since $\sigma$ is optimal for player I, for each real $b$, for $\mu_{\sigma,\tau_{b}}$-measure one many reals $a$, $G^{1}_{a}$ codes $S\cap\xi$ for some $\xi<\alpha$. Fix a real $b$.
By Corollary 18, every set of reals is $\mu_{\sigma,\tau_{b}}$-measurable and the ideal of $\mu_{\sigma,\tau_{b}}$-null sets is closed under well-ordered unions. Hence there is a unique $\xi_{b}<\alpha$ such that for $\mu_{\sigma,\tau_{b}}$-positive measure many reals $a$, $G^{1}_{a}$ codes $S\cap\xi_{b}$ and the set of reals $a$ such that $G^{1}_{a}$ codes $S\cap\xi$ for some $\xi<\xi_{b}$ is $\mu_{\sigma,\tau_{b}}$-measure zero. Let $C=\bigcup\{G^{1}_{a}\mid\text{$G^{1}_{a}$ codes $S\cap\xi_{b}$ for some $b$}\}$. Then for all reals $x$, $x$ is in $C$ if and only if there is a real $b$ such that for $\mu_{\sigma,\tau_{b}}$-positive measure many reals $a$, they code the same subset $S^{\prime}$ of $\gamma$, and no proper subsets of $S^{\prime}$ can be coded by $\mu_{\sigma,\tau_{b}}$-positive measure many reals, and $x\in G^{1}_{a}$ for some real $a$ such that $G^{1}_{a}$ codes $S^{\prime}$. Since $\Gamma$ is closed under $\exists^{\mathbb{R}}$ and $\forall^{\mathbb{R}}$, $C$ is in $\undertilde{\mathbf{\Gamma}}$. By induction hypothesis, for any $\xi<\alpha$, there is a real $b$ such that $G^{1}_{b}$ codes $S\cap\xi$. Since $\sigma$ is optimal for player I, $C$ codes $S\cap\alpha$, as desired.
@@
We show the claim by induction on $w$. Suppose it is done for all $v<w$. Then $A_{a^{*},w}$ codes $S\cap\xi$ where $\xi=\mathrm{sup}\{\eta(v)\mid v<w\}\geq\rho(w)$. Since $\tau$ is optimal for II, $C_{a^{*},w}$ codes $S\cap\eta$ for some $\eta>\xi$. Since $G^{1}_{g(w)}=C_{a^{*},w}$, setting $\eta(w)=\eta$, $\eta(w)>\rho(w)$ and $G^{1}_{g(w)}$ codes $S\cap\eta(w)$. ∎
@@ 4. $\mathsf{Bl}\text{-}\mathsf{AD}_{\mathbb{R}}$ and the Baire property
$\begin{array}[]{cccccc}\text{I}&s_{\pi_{0}}=s_{0}&&s_{\pi_{1}}&&\ldots\\ \text{II}&&s_{\pi_{0}+1},\ldots,s_{\pi_{0}-1}&&s_{\pi_{1}+1},\ldots,s_{\pi_{2}-1}$$\ldots\\ \end{array}$
@@ 5. $\mathsf{Bl}\text{-}\mathsf{AD}_{\mathbb{R}}$ and $\infty$-Borel codes
Take an arbitrary set $X$ and fix an injection $i_{X}\colon\text{OD}_{X}\to\text{HOD}_{X}$ which is ordinal definable with the parameter $X$. Then consider the Vopěnka algebra $\mathbb{P}_{\mathrm{V},X}$ in $\mathrm{HOD}_{X}$ as follows: $\mathbb{P}_{\mathrm{V},X}=\{i_{X}(A)\mid A\in\mathrm{OD}\text{ and }A\subseteq\wp(\omega)\}$. For $\in\mathbb{P}_{\mathrm{V},X}, pqp\leq q$ if $i_{X}^{-1}(p)\subseteq i_{X}^{-1}(q)$. It is easy to see that the properties of $\mathbb{P}_{\mathrm{V},X}$ do not depend on the choice of $i_{X}$, i.e., if there are two such injections, then the corresponding two preorders are isomorphic in $\mathrm{HOD}_{X}$. Vopěnka [23] proved that $\mathbb{P}_{\mathrm{V},\emptyset}$ is a complete Boolean algebra in HOD (when $X=\emptyset$) and each real in $V$ induces a $\mathbb{P}_{\mathrm{V},\emptyset}$-generic filter over HOD in the following way: For each real $x$ in $V$, the set $G_{x}=\{p\in\mathbb{P}_{\mathrm{V},\emptyset}\mid x\in i_{\emptyset}^{-1}(p)\}$ is a $\mathbb{P}_{\mathrm{V},\emptyset}$-generic filter over HOD and $x\in\text{HOD}[G_{x}]$. Conversely, if $G$ is a $\mathbb{P}_{\mathrm{V},\emptyset}$-generic filter over HOD, then the set $\bigcap\{i_{\emptyset}^{-1}(p)\mid p\in G\}$ is a singleton. We call the element of the singleton a Vopěnka real over HOD and denote it by $y_{G}$. Then $y_{G_{x}}=x$ for each real $x$ in $V$. The analogue of the above results holds for $\mathrm{HOD}_{X}$ for an arbitrary set $X$.
@@
Proof.
@@
The definition of strong $\infty$-Borel codes says that given $S,\gamma$ and $\pi$ as in the definition, for ‘stationary many’ $a$ in $\wp_{\omega_{1}}(\mathbb{R})$, $a$ is ‘good’ i.e., $B_{S\upharpoonright\pi[a]}\subseteq B_{S}$. Using the range-invariant determinacy (that follows from $\mathsf{Bl}\text{-}\mathsf{AD}_{\mathbb{R}}$ by Lemma 29), one can make it for ‘club many’ $a$ in $\wp_{\omega_{1}}(\mathbb{R})$, $a$ is ‘good’ in the same sense:
@@
We argue only for $\bigwedge_{n<\omega}S_{n}$ (the $\infty$-code $\bigvee_{n<\omega}S_{n}$ is similarly taken care of). Let $\pi\colon\mathbb{R}\to\gamma$ be a surjection. By Lemma 39, for each $n<\omega$, there is a function $f_{n}\colon\mathbb{R}^{<\omega}\to\mathbb{R}$ such that if $a$ is closed under $f_{n}$, then $S_{n}\upharpoonright\pi[a]$ is an $\infty$-Borel code and $B_{S_{n}\upharpoonright\pi[a]}\subseteq B_{S_{n}}$. Using $\mathsf{DC}$, we choose an $\omega$-sequence $(f_{n}\mid n<\omega)$ of such functions. Then one can find an $f\colon\mathbb{R}^{<\omega}\to\mathbb{R}$ such that if $a$ is closed under $f$, then so is each $f_{n}$. Then this $f$ witnesses that $\bigwedge_{n<\omega}S_{n}$ is a strong $\infty$-Borel code. ∎
@@
We now prove the key lemma for the theorem:

### v2

@@ Abstract.
The research in this paper was partially conducted when the authors stayed in the Mittag-Leffler Institute for the semester program “Mathematical Logic: Set Theory and Model Theory” in fall 2009. The authors are grateful to the Mittag-Leffler Institute for their hospitality and support. This paper contains results from the first author’s Ph.D. thesis [6]. The first author would like to thank Benedikt Löwe for his supervision. The first author’s research is partially supported by the Guangdong Philosophy and Social Science Foundation with grant number GD25YZX01. The authors would like to thank the referee for their helpful comments which improved the presentation of this paper.
@@ 1. Introduction
In 1998, Martin [16] proved that in most cases, Blackwell determinacy axioms follow from the corresponding determinacy axiom. Martin conjectured that they are equivalent, and many instances of equivalence have been shown (e.g., [17] and Martin’s proof of $\mathbf{\Pi}^{1}_{1}$ determinacy presented in [14, Corollary 3.9]). However, the general question, and in particular the most intriguing instance, namely whether $\mathsf{AD}$ and the axiom of Blackwell determinacy $\mathsf{Bl}\text{-}\mathsf{AD}$ are equivalent, remains open.
@@
Solovay [20] proved that the consistency of $\mathsf{ZF}$$\mathsf{DC}$$\mathsf{AD}_{\mathbb{R}}$ is strictly stronger than that of $\mathsf{ZF}$$\mathsf{AD}_{\mathbb{R}}$. Hence assuming $\mathsf{DC}$ is not optimal when considering the equivalence between $\mathsf{AD}_{\mathbb{R}}$ and $\mathsf{Bl}\text{-}\mathsf{AD}_{\mathbb{R}}$. While we do not know whether they are equivalent in $\mathsf{ZF}$$\mathsf{AC}_{\omega}(\mathbb{R})$, we show that they are equiconsistent:
@@ 2.1. Measure and category
Let $X$ be a set with at least two elements. The topology of $X^{\omega}$ is given by the product topology where each coordinate (i.e., $X$) is seen as the discrete space. We say $\mu$ is a Borel probability on $X^{\omega}$ if $\mu$ is a probability measure on the collection of Borel subsets of the space $X^{\omega}$.
@@
Suppose that every set of reals is Lebesgue measurable. Then for every Borel probability $\mu$ on the Cantor space $2^{\omega}$, every set of reals is $\mu$-measurable and the ideal of $\mu$-null sets is closed under well-ordered unions. Also, there is no injection from $\omega_{1}$ to the reals.
@@ 2.3. Blackwell games
For the arguments for the assertion that $\Gamma$ is a boldface pointclass as well as the closure of $\Gamma$ under countable unions and intersections, and $\exists^{\mathbb{R}}$, see [18, 2B.2]. For the closure of $\Gamma$ under $\forall^{\mathbb{R}}$, let $\Delta$ be the pointclass of all the Suslin and co-Suslin sets. Then by $\mathsf{Bl}\text{-}\mathsf{AD}$ and Theorem 13, every set of reals in $\Delta$ is determined. Now by the second periodicity theorem (see [18, 6C.3]), $\Gamma$ is closed under $\forall^{\mathbb{R}}$. ∎
@@ 2.6. Wadge games and Blackwell determinacy
where $\mu_{x,\tau}=\mu_{\sigma_{x},\tau}$ and $\sigma_{x}$ is the mixed strategy for player I concentrating on $x$ of measure one as in § 2.3. Notice that the right hand side of the equivalence is $\undertilde{\mathbf{\Sigma}}^{1}_{2}$ in $B$: in fact, the right hand side holds if and only if for all $n\in\omega$, there is a tree $T$ on $2$ such that $\mu_{x,\tau}(\{x\}\times[T])>1-\frac{1}{n+1}$ and for all $y\in[T]$, $y\in B$. (The statement $\mu_{x,\tau}(\{x\}\times[T])>1-\frac{1}{n+1}$’ is Borel in $x$ and $T$.) If player I has an optimal strategy in $\mathcal{G}_{\mathrm{W}}(A,B)$, then one can prove that $B$ is $\undertilde{\mathbf{\Sigma}}^{1}_{2}$ in $A^{\mathrm{c}}(=2^{\omega}\setminus A)$ in the same way and hence $B$ is $\undertilde{\mathbf{\Sigma}}^{1}_{2}$ in $A$. ∎
@@ 2.7. A weak version of Moschovakis’ Coding Lemma
Recall from § 2.3 that for a real $b$, $\tau_{b}$ is the mixed strategy for player II concentrating on $b$ of measure one. Since $\sigma$ is optimal for player I, for each real $b$, for $\mu_{\sigma,\tau_{b}}$-measure one many reals $a$, $G^{1}_{a}$ codes $S\cap\xi$ for some $\xi<\alpha$.
Fix a real $b^{\prime}$. By Corollary 18, every set of reals is $\mu_{\sigma,\tau_{b^{\prime}}}$-measurable and the ideal of $\mu_{\sigma,\tau_{b^{\prime}}}$-null sets is closed under well-ordered unions. Hence there is a unique $\xi_{b^{\prime}}<\alpha$ such that for $\mu_{\sigma,\tau_{b^{\prime}}}$-positive measure many reals $a$, $G^{1}_{a}$ codes $S\cap\xi_{b^{\prime}}$ and the set of reals $a$ such that $G^{1}_{a}$ codes $S\cap\xi$ for some $\xi<\xi_{b^{\prime}}$ is $\mu_{\sigma,\tau_{b^{\prime}}}$-measure zero.
Let $C=\bigcup\{G^{1}_{a}\mid\text{$G^{1}_{a}$ codes $S\cap\xi_{b}$ for some $b$}\}$. Then for all reals $x$, $x$ is in $C$ if and only if there is a real $b$ such that for $\mu_{\sigma,\tau_{b}}$-positive measure many reals $a$, they code the same subset $S^{\prime}$ of $\gamma$, and no proper subsets of $S^{\prime}$ can be coded by $\mu_{\sigma,\tau_{b}}$-positive measure many reals, and $x\in G^{1}_{a}$ for some real $a$ such that $G^{1}_{a}$ codes $S^{\prime}$. Since $\Gamma$ is closed under $\exists^{\mathbb{R}}$ and $\forall^{\mathbb{R}}$, $C$ is in $\undertilde{\mathbf{\Gamma}}$. By induction hypothesis, for any $\xi<\alpha$, there is a real $b$ such that $G^{1}_{b}$ codes $S\cap\xi$. Since $\sigma$ is optimal for player I, $C$ codes $S\cap\alpha$, as desired.
@@
We show the claim by induction on $w$. Fix a $w\in A$ with $\rho(w)<\alpha$. Suppose that for all $v<w$, the conclusion of the claim holds for $v$. We will verify the conclusion of the claim for $w$, i.e., there is some $\eta(w)<\alpha$ with $\rho(w)<\eta(w)$ such that $G^{1}_{g(w)}$ codes $S\cap\eta(w)$.
We first argue that $A_{a^{*},w}=\bigcup_{v<w}G^{1}_{g(v)}$. By the definition of $A_{a^{*},w}$, for all reals $x$, $x$ is in $A_{a^{*},w}$ if and only if for some $v<w$, $\{a^{*}\}(v)$ is defined and $\bigl(\{a^{*}\}(v),x\bigr)\in G^{1}$, i.e., $x$ is in $G^{1}_{\{a^{*}\}(v)}$. Since the function $(a,v^{\prime})\mapsto\pi^{\prime}(a,v^{\prime})$ is total and $\{a^{*}\}(v^{\prime})=\pi^{\prime}(a^{*},v^{\prime})$ for all reals $v^{\prime}$, $\{a^{*}\}(v)$ is defined for all $v<w$. Also, by the definition of $g$, $\{a^{*}\}(v)=g(v)$ for all $v<w$. Hence, for all reals $x$, $x$ is in $A_{a^{*},w}$ if and only if for some $v<w$, $x$ is in $G^{1}_{\{a^{*}\}(v)}=G^{1}_{g(v)}$. Therefore, $A_{a^{*},w}=\bigcup_{v<w}G^{1}_{g(v)}$, as desired.
By induction hypothesis, for all $v<w$, there is some $\eta(v)<\alpha$ with $\rho(v)<\eta(v)$ such that $G^{1}_{g(v)}$ codes $S\cap\eta(v)$. Since $A_{a^{*},w}=\bigcup_{v<w}G^{1}_{g(v)}$, letting $\xi=\mathrm{sup}\{\eta(v)\mid v<w\}$, $A_{a^{*},w}$ codes $S\cap\xi$. Also, since $\eta(v)>\rho(v)$ for all $v<w$, $\xi=\mathrm{sup}\{\eta(v)\mid v<w\}\geq\mathrm{sup}\{\rho(v)+1\mid v<w\}=\rho(w)$.
We next argue that $C_{a^{*},w}$ codes $S\cap\eta$ for some $\eta>\xi$. By the definition of $C_{a^{*},w}$, for some subset $S^{\prime}$ of $\gamma$, a real $x$ is in $C_{a^{*},w}$ if and only if for $\mu_{\sigma_{\pi(a^{*},w)},\tau}$-positive measure many reals $b$, $G^{1}_{b}$ codes $S^{\prime}$, no proper subsets of $S^{\prime}$ can be coded by $\mu_{\sigma_{\pi(a^{*},w)},\tau}$-positive measure many reals, and $x$ is in $G^{1}_{b}$ for some $b$ such that $G^{1}_{b}$ codes $S^{\prime}$. Hence $C_{a^{*},w}$ also codes this $S^{\prime}$. By the definition of $\pi$, $A_{a^{*},w}=G^{1}_{\pi(a^{*},w)}$. Since $A_{a^{*},w}=G^{1}_{\pi(a^{*},w)}$ codes $S\cap\xi$ and $\tau$ is an optimal strategy for player II, the set $S^{\prime}$ must be of the form $S\cap\eta$ for some $\eta>\xi$. Hence $C_{a^{*},w}$ codes $S\cap\eta$ for some $\eta>\xi$, as desired.
Set $\eta(w)=\eta$. Then since $\eta>\xi\geq\rho(w)$, $\eta(w)>\rho(w)$. By the definition of $g$, $g(w)=\{a^{*}\}(w)=\pi^{\prime}(a^{*},w)$. So $G^{1}_{g(w)}=G^{1}_{\pi^{\prime}(a^{*},w)}=C_{a^{*},w}$, where the second equality follows from the property of $\pi^{\prime}$. Since $C_{a^{*},w}$ codes $S\cap\eta=S\cap\eta(w)$, $G^{1}_{g(w)}$ codes $S\cap\eta(w)$, as desired. ∎
@@ 4. $\mathsf{Bl}\text{-}\mathsf{AD}_{\mathbb{R}}$ and the Baire property
$\begin{array}[]{cccccc}\text{I}&s_{\pi_{0}}=s_{0}&&s_{\pi_{1}}&&\ldots\\ \text{II}&&s_{\pi_{0}+1},\ldots,s_{\pi_{1}-1}&&s_{\pi_{1}+1},\ldots,s_{\pi_{2}-1}$$\ldots\\ \end{array}$
@@ 5. $\mathsf{Bl}\text{-}\mathsf{AD}_{\mathbb{R}}$ and $\infty$-Borel codes
Take an arbitrary set $X$ and fix an injection $i_{X}\colon\text{OD}_{X}\to\text{HOD}_{X}$ which is ordinal definable with the parameter $X$. Then consider the Vopěnka algebra $\mathbb{P}_{\mathrm{V},X}$ in $\mathrm{HOD}_{X}$ as follows: $\mathbb{P}_{\mathrm{V},X}=\{i_{X}(A)\mid A\in\mathrm{OD}\text{ and }A\subseteq\wp(\omega)\}$. For $\in\mathbb{P}_{\mathrm{V},X}, pqp\leq q$ if $i_{X}^{-1}(p)\subseteq i_{X}^{-1}(q)$. It is easy to see that the properties of $\mathbb{P}_{\mathrm{V},X}$ do not depend on the choice of $i_{X}$, i.e., if there are two such injections, then the corresponding two preorders are isomorphic in $\mathrm{HOD}_{X}$. Vopěnka [23] proved that $\mathbb{P}_{\mathrm{V},\emptyset}$ is a complete Boolean algebra in HOD (when $X=\emptyset$) and each real in $V$ induces a $\mathbb{P}_{\mathrm{V},\emptyset}$-generic filter over HOD in the following way: For each real $x$ in $V$, the set $G_{x}=\{p\in\mathbb{P}_{\mathrm{V},\emptyset}\mid x\in i_{\emptyset}^{-1}(p)\}$ is a $\mathbb{P}_{\mathrm{V},\emptyset}$-generic filter over HOD and $x\in\text{HOD}[G_{x}]$. [^4] Here is the argument for $x\in\text{HOD}[G_{x}]$: For each $n\in\omega$ and $i=0,1$, let $N_{n,i}=\{y\in 2^{\omega}\mid y(n)=i\}$. Then the function $(n,i)\mapsto i_{\emptyset}(N_{n,i})$ is OD and each $i_{\emptyset}(N_{n,i})$ is in HOD. So the function $(n,i)\mapsto i_{\emptyset}(N_{n,i})$ is in HOD. Now for all $n$ and $i$, $x(n)=i\iff x\in N_{n,i}\iff i_{\emptyset}(N_{n,i})\in G_{x}$. Hence $x$ can be computed from $G_{x}$ and the function $(n,i)\mapsto i_{\emptyset}(N_{n,i})$. Therefore, $x\in\text{HOD}[G_{x}]$. Conversely, if $G$ is a $\mathbb{P}_{\mathrm{V},\emptyset}$-generic filter over HOD, then the set $\bigcap\{i_{\emptyset}^{-1}(p)\mid p\in G\}$ is a singleton. We call the element of the singleton a Vopěnka real over HOD and denote it by $y_{G}$. Then $y_{G_{x}}=x$ for each real $x$ in $V$. The analogue of the above results holds for $\mathrm{HOD}_{X}$ for an arbitrary set $X$.
@@
Therefore, a given $\infty$-Borel code $S$ is strong if and only if there are ‘stationarily many’ $a$ in $\wp_{\omega_{1}}(\mathbb{R})$ such that for any real $x$, if $\mathrm{L}_{\overline{\gamma}}[\overline{S},x]\vDash\phi(\overline{S},x)$, then $\mathrm{L}_{\gamma}[S,x]\vDash\phi(S,x)$, where $\gamma$, $\overline{\gamma}$, and $\overline{S}$ are as in Theorem 37.
Proof of Theorem 37.
@@
The definition of strong $\infty$-Borel codes says that given $S,\gamma$ and $\pi$ as in the definition, for ‘stationarily many’ $a$ in $\wp_{\omega_{1}}(\mathbb{R})$, $a$ is ‘good’ i.e., $B_{S\upharpoonright\pi[a]}\subseteq B_{S}$. Using the range-invariant determinacy (that follows from $\mathsf{Bl}\text{-}\mathsf{AD}_{\mathbb{R}}$ by Lemma 29), one can make it for ‘club many’ $a$ in $\wp_{\omega_{1}}(\mathbb{R})$, $a$ is ‘good’ in the same sense:
@@
We argue only for $\bigwedge_{n<\omega}S_{n}$ (the $\infty$-code $\bigvee_{n<\omega}S_{n}$ is similarly taken care of). Let $\pi\colon\mathbb{R}\to\gamma$ be a surjection. By Lemma 39, for each $n<\omega$, there is a function $f_{n}\colon\mathbb{R}^{<\omega}\to\mathbb{R}$ such that if $a$ is closed under $f_{n}$, then $S_{n}\upharpoonright\pi[a]$ is an $\infty$-Borel code and $B_{S_{n}\upharpoonright\pi[a]}\subseteq B_{S_{n}}$. Using $\mathsf{DC}$, we choose an $\omega$-sequence $(f_{n}\mid n<\omega)$ of such functions. Then one can find an $f\colon\mathbb{R}^{<\omega}\to\mathbb{R}$ such that if $a$ is closed under $f$, then $a$ is closed under each $f_{n}$. Then this $f$ witnesses that $\bigwedge_{n<\omega}S_{n}$ is a strong $\infty$-Borel code. ∎
@@
Here is the key lemma for the theorem: